Question 1
If we mixed 45cm^3 of 2m NaOH with 20cm^3 of 1m HCL and obtained a temperature difference of 9.4°C, what would be the heat of reaction, the enthalpy?.
Question 2
If we added 10cm^3 HCL of an unknown concentration to 30cm^3 of 2M NaOH and obtained a temperature change of 4.5°C and then added 20cm^3 of the same unknown acid to 30cm^3 of 2M NaOH and obtained a temperature change of 9.0°C, what is the concentration of the acid? Use the enthalpy value from question 1
1
Expert's answer
2019-05-27T06:10:32-0400

Solution.

Question 1.


"Q =m \times c \times \Delta T"

"m = m(HCl) + m(NaOH)"

"Q = 0.00431*4200*9.4 = 170.1588 J"

"\Delta H = -170.1588 J"

Question 2.


"m = \frac{Q}{c \times \Delta T}"

m = 4.5 g


"m(NaOH) = 0.03*3*40 = 3.6"

m(HCl) = 0.9 g


"c(HCl) = \frac{0.9}{35.5*0.01} = 1.26 M"

Answer:

Question 1

Q = 170.1588

H = -170.1588

Question 2.

c(HCl) = 1.26 M